OpenStudy (anonymous):

Solve the compound inequality. Graph the solution. –2 ≤ 2x – 4 < 4

OpenStudy (texaschic101):

-2 <= 2x - 4 < 4 ---- add 4 4 - 2 <= 2x < 4 + 4 2 <= 2x < 8 --- divide by 2 2/2 <= x < 8/2 1 <= x < 4 so we have x is greater or equal to 1 and less then 4 closed circle on 1, open circle on 4, shading in between

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No, I think I've got it! But thank you so much :) I really appreciate it!

OpenStudy (texaschic101):

no problem :)